# A review of Bayesian sensor-based estimation and uncertainty quantification of aerodynamic flows

### v2

@@ 4 Neural networks and other learned mappings
with $\nabla_{\mathsf{y}}\tilde{g}$ denoting the Jacobian matrix of the operator $\tilde{g}$, which is evaluated at the expected value (mean) of a measurement. While this method provides a computationally efficient uncertainty quantification strategy, it is inherently limited in highly nonlinear networks, where local linearization fails to capture complex input-output mappings. To address this, Bibi et al. [37] derived analytical, layer-wise expressions for propagating the first and second moments of input noise when the noise follows a Gaussian distribution. It should also be noted that this local linearization is the same idea behind the state-space Gramian (24) defined on the observation operator, and this can be easily extended to the inverse operator. The Gramian definition mitigates the effect of nonlinearity by computing expectations over the distribution of inputs, rather than only evaluating the Jacobian at mean.
We could alternatively pursue an approach that combines the training and uncertainty quantification into a single task [35]. To understand how this works, it is important to remember that training a mathematical model for least-squares regression can be viewed as finding a normal distribution $\mathcal{N}(\tilde{g}(\mathsf{y};w),\Sigma)$ that maximizes the likelihood over the set of (random) training outputs [2]. We perform this maximization with respect to the weights $w$ of the network and the (as yet unknown) variance $\Sigma$ of the outputs. After training, the newly-determined $\Sigma$ is called the aleatoric (or data) uncertainty. When the inputs and outputs are both free of noise, as in the training of the deterministic network $\tilde{g}(\mathsf{y};w)$, then this perspective of regression is not particularly helpful. However, suppose we start with the same training data $D$, but during each training epoch we add measurement noise drawn from $\mathcal{N}(0,\Sigma_{\mathcal{E}})$ “on the fly” to the inputs. We do not explicitly add noise to the outputs $\boldsymbol{\xi}_{j}$, but by virtue of their pairing with random inputs we can also view them as random, as well. Training the network with this noisy data allows us to determine the aleatoric uncertainty of the outputs, and this uncertainty is attributable to the noisy inputs.
In fact, we can embrace this perspective further by expanding the network to predict both the mean and the variance as functions of the input, as shown in Fig. 15, adapted from [35]. Such a neural network is called heteroscedastic [38], since the variance is not global, but dependent on the input. In our example, this input is pressure measurements, and the architecture illustrated in Fig. 15 is designed to learn the parameters of a multivariate Gaussian distribution in the latent space,
@@
The network is trained with the objective function defined as the negative log of the likelihood of the output training data, assuming this distribution. The aleatoric uncertainty quantification in the predicted latent vector, arising from noise in surface pressure measurements, has been thoroughly investigated in [35] using such a heteroscedastic network. An advantage of such a model is that it provides a generative model for the latent vectors for further Monte Carlo tasks. In particular, for a given measurement vector $\mathsf{y}$, samples can be drawn from the normal distribution (38) and then propagated through the decoder and lift networks to quantify uncertainty in the flow field and the lift, respectively.
The methods discussed above represent common approaches for propagating data noise through NNs, thereby accounting for the data (or aleatoric) uncertainty in predictions. In these formulations, we assume that the mapping $\boldsymbol{\xi}=\tilde{g}(\mathsf{y};w)$ is learned perfectly, meaning the NN parameters $w$ accurately capture the underlying true physical function. However, in practice, NNs often approximate complex dynamics with inherent modeling errors. These discrepancies arise due to limitations in the learned function, structural simplifications, or insufficient training data, leading to epistemic (or model) uncertainty–—the uncertainty associated with the model itself. Model uncertainty is particularly crucial in domains such as safety-critical control systems [39], medical diagnosis [40], and autonomous driving [41], where reliable uncertainty quantification is essential for risk-aware decision-making.
Taking the same inverse mapping $\boldsymbol{\xi}=\tilde{g}(\mathsf{y};w)$ as discussed in the preceding paragraphs, the uncertainty in the model parameters $w$, denoted as $\pi(w|D)$, can be estimated using various approaches. One widely used method is Deep Ensemble, where a deterministic neural network is trained independently $N$ times, producing a set of models $\tilde{g}^{i}$, for $i=1,\cdots,N$. The predicted outputs from these independently trained models, $\{\boldsymbol{\xi}^{i}\}=\{\tilde{g}^{i}(\mathsf{y};w)\}_{i=1}^{N}$ can be interpreted as samples from the predictive distribution, providing an empirical estimation of epistemic uncertainty. This approach has been extensively used to enhance the robustness of neural networks [42, 43]. However, its major limitation is computational inefficiency, as it requires training multiple networks independently—an impractical approach for large-scale or deep architectures where training costs are prohibitively high.
@@
The intuition behind Eq. (39) lays the foundation for a transformative concept in the field of computer science: the Bayesian Neural Network (BNN). BNNs [44, 45, 46] are stochastic NNs trained using Bayesian inference. In a BNN, the prior distribution over parameters can either be specified based on prior knowledge (such as from a pre-trained deterministic network [47]) or learned during training. Given access to the posterior distribution, $\pi(w|D)$, one can sample from it to approximate the predictive distribution, enabling principled uncertainty quantification and robust decision-making. This predictive distribution is computed as
@@
However, obtaining the exact posterior distribution is challenging due to the intractability of the denominator in Eq. (39). Instead of explicitly computing this integral, MCMC methods [48] provide a means of sampling from the true posterior distribution by leveraging the proportionality relationship: $\pi(w|D)\propto\pi(D|w)\ \pi(w)$. While MCMC is theoretically exact, it becomes computationally prohibitive for high-dimensional parameter spaces, particularly in deep NNs.
As a more scalable alternative, variational inference (VI) [49] approximates the true posterior $\pi(w|D)$ with a tractable probability distribution $q_{\phi}(w)$, parameterized by $\phi$. The network is trained to optimize these parameters so that $q_{\phi}(w)$ closely resembles the true posterior. The quality of this approximation is measured using the Kullback-Leibler (KL) divergence [6], which quantifies the discrepancy between the variational distribution and the true posterior:
@@
Despite the effectiveness of VI in BNNs, directly learning the parameters of the approximate posterior remains computationally expensive and challenging to converge. To address this, Monte Carlo (MC) dropout, introduced by Gal and Ghahramani [50, 51], provides an efficient approximation of the posterior distribution using dropout as a form of variational inference. They mathematically demonstrated that placing a random dropout layer after each dense layer is equivalent to performing VI in a deep Gaussian process. MC dropout is easy to implement and has been utilized by Mousavi and Eldredge [35] to quantify the epistemic uncertainty in disturbed aerodynamics.
@@ 5 Conclusions and future directions
We have only discussed a way of thinking about uncertainty in these learned operators, and there is still much more to do to exploit them in the sequential estimation framework. In this work, we have assumed that an autoencoder is trained from snapshots of the full flow field, free from noise, and we have treated the resulting latent vectors as a perfect surrogate of the full flow. Suppose, instead, we train an autoencoder from snapshots of planar PIV data in a three-dimensional flow past a wing. The latent space identified by this is inherently noisy and missing out-of-plane flow information. A variational autoencoder [52] would serve this task well. We might use the Bayesian framework to train an observation operator to assimilate both PIV data and measurements from pressure sensors distributed along both the span and chord of the wing, and exploit our knowledge of Navier–Stokes to fill in the remaining details of the three-dimensional flow around the wing.
@@
There are a number of other interesting and important aspects of flow estimation that require further investigation. We have not addressed the important issue of sensor placement, though the probabilistic framework certainly provides a powerful toolbox for addressing this. The informative subspaces we have discussed here are obviously germane to sensors’ observability over time. This review has not devoted any explicit consideration of the estimation of turbulent flows and their wide spectrum of length and time scales. It is important to note that some of the biggest advances to data assimilation of flows have been made in the area of numerical weather prediction, in which the breadth of these spectra is as large as any we encounter. Sparse sensors record the footprints of the largest scales of weather systems, and our knowledge of the governing equations of these flows—or, at least, our characterization of their statistics—enables us to rationalize these sensor measurements with wider knowledge of the flow structure. In linear stochastic estimation (LSE) [53, 54], developed by Adrian and co-workers, the mean velocity field in a turbulent flow is conditioned on events, i.e., pre-chosen characteristics of the velocity field at certain locations and times. The resulting “conditional eddy” structure is determined by the two-point correlations of the turbulent flow and provides a mean of inferring the expected behavior in the vicinity of certain observations. It should also added that the dynamics of turbulent flows have a much sharper sensitivity to the state than unsteady low Reynolds numbers. Data assimilation provides a means of mitigating this tendency toward spatio-temporal chaos. The recent review by Zaki [55] provides an excellent discussion of this view and its mathematical underpinnings.
@@
Flow and flight control are implicit motivations for much of what we have discussed in this paper, but we have not addressed the relationship between the estimated flow state and the actions that we take to achieve some control objective. In reinforcement learning, a machine learning framework that leverages experiential interactions with the environment to learn a control strategy, this strategy is typically expressed as a conditional probability, $\pi(\mathsf{a}|\mathsf{x})$—that is, given a system state, what is the distribution of actions $\mathsf{a}$ that we could take to maximize the expected rewards over time. One could reasonably ask, if we can successfully estimate $\mathsf{x}$ from observations $\mathsf{y}$, then could we condition the actions on $\mathsf{y}$ alone? To what extent do we explicitly need to estimate the state? These questions are critical for efficient training of successful reinforcement learning policies in fluid dynamics [56].
@@ Appendix A Derivation of posterior mean and covariance in linear equations
$-\frac{1}{2}\begin{pmatrix}\mathsf{x}-\mu_{\mathsf{x}}\\ \mathsf{y}^{\ast}-\mu_{\mathsf{y}}\end{pmatrix}^{T}\begin{bmatrix}\Sigma_{\mathsf{x}\mathsf{x}}&\Sigma_{\mathsf{x}\mathsf{y}}\\ \Sigma_{\mathsf{y}\mathsf{x}}&\Sigma_{\mathsf{y}\mathsf{y}}\end{bmatrix}^{-1}\begin{pmatrix}\mathsf{x}-\mu_{\mathsf{x}}\\ \mathsf{y}^{\ast}-\mu_{\mathsf{y}}\end{pmatrix}$ (48)
into a quadratic form for $\mathsf{x}$ only, treating $\mathsf{y}^{\ast}$ as a constant [2]. The result is
$-\frac{1}{2}(\mathsf{x}-\mu_{\mathsf{x}|\mathsf{y}^{\ast}})^{T}\Sigma_{\mathsf{x}|\mathsf{y}^{\ast}}^{-1}(\mathsf{x}-\mu_{\mathsf{x}|\mathsf{y}^{\ast}}),$ (49)
@@ Appendix B Ensemble approximations of mean and covariances
An important note about the ensemble covariance matrices (60)–(62): their rank is, at most, $M-1$. Why? Each term in the sums is a rank-1 matrix, since it projects a vector onto a space spanned by a single column vector, e.g., $\mathsf{x}^{i}-\hat{\mu}_{\mathsf{x}}$. One might be tempted to think that the overall rank of the matrix is therefore $M$ (a sum of $M$ rank-1 matrices), but it is actually $M-1$, because these column vectors are not independent: they sum to zero, by virtue of the definitions of the means (58) and (59).

### v3

@@ 4 Neural networks and other learned mappings
with $\nabla_{\mathsf{y}}\tilde{g}$ denoting the Jacobian matrix of the operator $\tilde{g}$, which is evaluated at the expected value (mean) of a measurement. While this method provides a computationally efficient uncertainty quantification strategy, it is inherently limited in highly nonlinear networks, where local linearization fails to capture complex input-output mappings. To address this, Bibi et al. [bibi2018analytic] derived analytical, layer-wise expressions for propagating the first and second moments of input noise when the noise follows a Gaussian distribution. It should also be noted that this local linearization is the same idea behind the state-space Gramian (24) defined on the observation operator, and this can be easily extended to the inverse operator. The Gramian definition mitigates the effect of nonlinearity by computing expectations over the distribution of inputs, rather than only evaluating the Jacobian at the mean.
We could alternatively pursue an approach that combines the training and uncertainty quantification into a single task [mousavi2025loworderflowreconstructionuncertainty]. To understand how this works, it is important to remember that training a mathematical model for least-squares regression can be viewed as finding a normal distribution $\mathcal{N}(\tilde{g}(\mathsf{y};w),\Sigma)$ that maximizes the likelihood over the set of (random) training outputs [bishop2006pattern]. We perform this maximization with respect to the weights $w$ of the network and the (as yet unknown) variance $\Sigma$ of the outputs. After training, the newly-determined $\Sigma$ is called the aleatoric (or data) uncertainty. When the inputs and outputs are both free of noise, as in the training of the deterministic network $\tilde{g}(\mathsf{y};w)$, then this perspective of regression is not particularly helpful. However, suppose we start with the same training data $D$, but during each training epoch, to the inputs we add measurement noise drawn from $\mathcal{N}(0,\Sigma_{\mathcal{E}})$ “on the fly”. We do not explicitly add noise to the outputs $\boldsymbol{\xi}_{j}$, but by virtue of their pairing with random inputs we can also view them as random, as well. Training the network with this noisy data allows us to determine the aleatoric uncertainty of the outputs, and this uncertainty is attributable to the noisy inputs.
In fact, we can embrace this perspective further by expanding the network to predict both the mean and the variance as functions of the input, as shown in Fig. 15, adapted from [mousavi2025loworderflowreconstructionuncertainty]. Such a neural network is called heteroscedastic [kendall2017uncertainties], since the variance is not global, but dependent on the input. In our example, this input is pressure measurements, and the architecture illustrated in Fig. 15 is designed to learn the parameters of a multivariate Gaussian distribution in the latent space,
@@
Measurement noise is translated to the predicted latent vector by learning the parameters of this distribution. The right panel in Fig. 15 depicts the low-dimensional latent space representation of disturbed flow fields.
The network is trained with the objective function defined as the negative log of the likelihood of the output training data, assuming this distribution. The aleatoric uncertainty quantification in the predicted latent vector, arising from noise in surface pressure measurements, has been thoroughly investigated in [mousavi2025loworderflowreconstructionuncertainty] using such a heteroscedastic network. An advantage of such a model is that it provides a generative model for the latent vectors for further Monte Carlo tasks. In particular, for a given measurement vector $\mathsf{y}$, samples can be drawn from the normal distribution (38) and then propagated through the decoder and lift networks to quantify uncertainty in the flow field and the lift, respectively.
The methods discussed above represent common approaches for propagating data noise through NNs, thereby accounting for the data (or aleatoric) uncertainty in predictions. In these formulations, we assume that the mapping $\boldsymbol{\xi}=\tilde{g}(\mathsf{y};w)$ is learned perfectly, meaning the NN parameters $w$ accurately capture the underlying true physical function. However, in practice, NNs often approximate complex dynamics with inherent modeling errors. These discrepancies arise due to limitations in the learned function, structural simplifications, or insufficient training data, leading to epistemic (or model) uncertainty–—the uncertainty associated with the model itself. Model uncertainty is particularly crucial in domains such as safety-critical control systems [le2018uncertainty], medical diagnosis [laves2019quantifying], and autonomous driving [shafaei2018uncertainty], where reliable uncertainty quantification is essential for risk-aware decision-making.
Taking the same inverse mapping $\boldsymbol{\xi}=\tilde{g}(\mathsf{y};w)$ as discussed in the preceding paragraphs, the uncertainty in the model parameters $w$, denoted as $\pi(w|D)$, can be estimated using various approaches. One widely used method is Deep Ensemble, where a deterministic neural network is trained independently $N$ times, producing a set of models $\tilde{g}^{i}$, for $i=1,\cdots,N$. The predicted outputs from these independently trained models, $\{\boldsymbol{\xi}^{i}\}=\{\tilde{g}^{i}(\mathsf{y};w)\}_{i=1}^{N}$ can be interpreted as samples from the predictive distribution, providing an empirical estimation of epistemic uncertainty. This approach has been extensively used to enhance the robustness of neural networks [ren2020deep, mcdermott2019deep]. However, its major limitation is computational inefficiency, as it requires training multiple networks independently—an impractical approach for large-scale or deep architectures where training costs are prohibitively high.
@@
The intuition behind Eq. (39) lays the foundation for a transformative concept in the field of computer science: the Bayesian Neural Network (BNN). BNNs [lampinen2001bayesian, titterington2004bayesian, goan2020bayesian] are stochastic NNs trained using Bayesian inference. In a BNN, the prior distribution over parameters can either be specified based on prior knowledge (such as from a pre-trained deterministic network [krishnan2019efficient]) or learned during training. Given access to the posterior distribution, $\pi(w|D)$, one can sample from it to approximate the predictive distribution, enabling principled uncertainty quantification and robust decision-making. This predictive distribution is computed as
@@
However, obtaining the exact posterior distribution is challenging due to the intractability of the denominator in Eq. (39). Instead of explicitly computing this integral, MCMC methods [bardenet2017markov] provide a means of sampling from the true posterior distribution by leveraging the proportionality relationship: $\pi(w|D)\propto\pi(D|w)\ \pi(w)$. While MCMC is theoretically exact, it becomes computationally prohibitive for high-dimensional parameter spaces, particularly in deep NNs.
As a more scalable alternative, variational inference (VI) [blei2017variational] approximates the true posterior $\pi(w|D)$ with a tractable probability distribution $q_{\phi}(w)$, parameterized by $\phi$. The network is trained to optimize these parameters so that $q_{\phi}(w)$ closely resembles the true posterior. The quality of this approximation is measured using the Kullback-Leibler (KL) divergence [kullback1951information], which quantifies the discrepancy between the variational distribution and the true posterior:
@@
Despite the effectiveness of VI in BNNs, directly learning the parameters of the approximate posterior remains computationally expensive and challenging to converge. To address this, Monte Carlo (MC) dropout, introduced by Gal and Ghahramani [gal2016dropoutbayesianapproximationappendix, gal2016dropout], provides an efficient approximation of the posterior distribution using dropout as a form of variational inference. They mathematically demonstrated that placing a random dropout layer after each dense layer is equivalent to performing VI in a deep Gaussian process. MC dropout is easy to implement and has been utilized by Mousavi and Eldredge [mousavi2025loworderflowreconstructionuncertainty] to quantify the epistemic uncertainty in disturbed aerodynamics.
@@ 5 Conclusions and future directions
We have only discussed a way of thinking about uncertainty in these learned operators, and there is still much more to do to exploit them in the sequential estimation framework. In this work, we have assumed that an autoencoder is trained from snapshots of the full flow field, free from noise, and we have treated the resulting latent vectors as a perfect surrogate of the full flow. Suppose, instead, we train an autoencoder from snapshots of planar PIV data in a three-dimensional flow past a wing. The latent space identified by this is inherently noisy and missing out-of-plane flow information. A variational autoencoder [kingma2019introduction] would serve this task well. We might use the Bayesian framework to train an observation operator to assimilate both PIV data and measurements from pressure sensors distributed along both the span and chord of the wing, and exploit our knowledge of Navier–Stokes to fill in the remaining details of the three-dimensional flow around the wing.
@@
There are a number of other interesting and important aspects of flow estimation that require further investigation. We have not addressed the important issue of sensor placement, though the probabilistic framework certainly provides a powerful toolbox for addressing this. The informative subspaces we have discussed here are obviously germane to sensors’ observability over time. This review has not devoted any explicit consideration of the estimation of turbulent flows and their wide spectrum of length and time scales. It is important to note that some of the biggest advances to data assimilation of flows have been made in the area of numerical weather prediction, in which the breadth of these spectra is as large as any we encounter. Sparse sensors record the footprints of the largest scales of weather systems, and our knowledge of the governing equations of these flows—or, at least, our characterization of their statistics—enables us to rationalize these sensor measurements with wider knowledge of the flow structure. In linear stochastic estimation (LSE) [adrian1975role, adrian1988stochastic], developed by Adrian and co-workers, the mean velocity field in a turbulent flow is conditioned on events, i.e., pre-chosen characteristics of the velocity field at certain locations and times. The resulting “conditional eddy” structure is determined by the two-point correlations of the turbulent flow and provides a mean of inferring the expected behavior in the vicinity of certain observations. It should also added that the dynamics of turbulent flows have a much sharper sensitivity to the state than unsteady low Reynolds numbers. Data assimilation provides a means of mitigating this tendency toward spatio-temporal chaos. The recent review by Zaki [zaki2024turbulence] provides an excellent discussion of this view and its mathematical underpinnings.
@@
Flow and flight control are implicit motivations for much of what we have discussed in this paper, but we have not addressed the relationship between the estimated flow state and the actions that we take to achieve some control objective. In reinforcement learning, a machine learning framework that leverages experiential interactions with the environment to learn a control strategy, this strategy is typically expressed as a conditional probability, $\pi(\mathsf{a}|\mathsf{x})$—that is, given a system state, what is the distribution of actions $\mathsf{a}$ that we could take to maximize the expected rewards over time. One could reasonably ask, if we can successfully estimate $\mathsf{x}$ from observations $\mathsf{y}$, then could we condition the actions on $\mathsf{y}$ alone? To what extent do we explicitly need to estimate the state? These questions are critical for efficient training of successful reinforcement learning policies in fluid dynamics [beckers2024deep].
@@ Appendix A Derivation of posterior mean and covariance in linear equations
$\frac{1}{2}\begin{pmatrix}\mathsf{x}-\mu_{\mathsf{x}}\\ \mathsf{y}^{\ast}-\mu_{\mathsf{y}}\end{pmatrix}^{T}\begin{bmatrix}\Sigma_{\mathsf{x}\mathsf{x}}&\Sigma_{\mathsf{x}\mathsf{y}}\\ \Sigma_{\mathsf{y}\mathsf{x}}&\Sigma_{\mathsf{y}\mathsf{y}}\end{bmatrix}^{-1}\begin{pmatrix}\mathsf{x}-\mu_{\mathsf{x}}\\ \mathsf{y}^{\ast}-\mu_{\mathsf{y}}\end{pmatrix}$ (48)
into a quadratic form for $\mathsf{x}$ only, treating $\mathsf{y}^{\ast}$ as a constant [bishop2006pattern]. The result is
$\frac{1}{2}(\mathsf{x}-\mu_{\mathsf{x}|\mathsf{y}^{\ast}})^{T}\Sigma_{\mathsf{x}|\mathsf{y}^{\ast}}^{-1}(\mathsf{x}-\mu_{\mathsf{x}|\mathsf{y}^{\ast}}),$ (49)
@@ Appendix B Ensemble approximations of mean and covariances
An important note about the ensemble covariance matrices (60)–(62): their rank is, at most, M1M-1. Why? Each term in the sums is a rank-1 matrix, since it projects a vector onto a space spanned by a single column vector, e.g., $\mathsf{x}^{i}-\hat{\mu}_{\mathsf{x}}$. One might be tempted to think that the overall rank of the matrix is therefore $M$ (a sum of $M$ rank-1 matrices), but it is actually M1M-1, because these column vectors are not independent: they sum to zero, by virtue of the definitions of the means (58) and (59).